. We are in a world of no corporate taxes. Markets in finance and investments are efficient. The risk-free rate of interest is 2.5% and the expected equity premium is 4%. In the competitive market for Electrical Equipment, all companies operate extremely efficiently. One such company is Safe Electrics (SE). They are currently an all equity company. Suppose the firm is currently valued at $10,000,000, which is the value of its operating assets. It has an expected return on operating assets of 6%. It doesnt pay any dividends. Assume there are no corporate taxes (in part A and B). (a) What is the expected earnings of SE during the next year (year 1)? If there are 500,000 shares in SE then what is the expected price per share at the end of year 1? (10 marks) (b) Suppose an investor wishes to generate a higher expected return than that offered by SE. Demonstrate how the investor could increase their expected return to 7.75% by borrowing money themselves. Illustrate for an investor who (at end of year 1) holds 1,000 shares in SE and has no borrowing. What level of homemade leverage is required for them to do this? (20 marks) (c) In a world with corporate taxes, firms should aim to take on as much debt as possible in order to maximise the value of tax shields. To what extent do you agree with this statement? (20 marks) Total: 50 marks)
Answer #1

(a) Since There are no corporate taxes, entire return earned by the company can be distributed among the equity share holders.

given risk free interest rate : 2.5%

Expected Equity premium: 4%

Expected return on operating assets: 6%

Total Value of operating assets: 10,000,000

Expected earnings = (Net Operating assets) * (Expected Return on Operating assets)

Expected earnings = 10,000,000*6% = $ 600,000

If there are 500,000 equity shares in Company SE

Expected Price per share at end of year 1 = (Net operating assets at the end of year 1)/ No of Equity Shares

Expected Price per share at end of year 1 = 10,600,000/500,000 = $ 21.20/share

(b) If an Investor is Holding 1000 shares

Present Value of his holding = 1000*20 = $ 20000

Suppose he purchased another 500 shares by Borrowing $ 10000 from outside @ 2.50% risk free Interest

His total Investment in SE company shares = $ 30000

Total earnings at the end of year 1 (A) = 30000*6% = $ 1800

Interest to be paid to Lender (B) = 20000*2.5% = $ 250

Net earnings by the investor = $ 1550

Return on his investment = 1550/20000 = 7.75%

Therefore if investor borrows $10000 from outside he can earn expected return on investment 7.75%


(c) Following Illustration effect of Corporate tax on Companies capital structure

Total Equity firm

Half equity

Half Debt Firm

Total Debt Firm
Equity $ 10,000 $ 5,000 0
Debt 0 $ 5,000 $ 10,000
Capital $ 10,000 $ 10,000 $ 10,000
EBIT (10%) $ 1,000 $ 1,000 $ 1,000
Interest (5%) 0 ($ 250) ($ 500)
EBT $ 1,000 $ 750 $ 500
TAX (50%) $ 500 $ 375 $ 250
EAT $ 500 $ 375 $ 250
ROE (EAT/EQUITY) 5% 7.5% > 7.5%

Strongly disagree:

Rationale: A firm should establish its capital structure in a prudent way.

Financial leverage only works well only when assets purchased through the borrowed funds earn more than Cost of debt. Otherwise it becomes very difficult to operate in a competitive environment.

The level of dependence on external borrowing is decided based on the industry in which company operates.

As shown above all debt company earnings becomes less but ROE becomes more than other two companies.